# A Real-time Multimodal Transformer Neural Network-powered Wildfire Forecasting System

### v1

@@ Abstract
Due to climate change, the extreme wildfire has become one of the most dangerous natural hazards to human civilization. Even though, some wildfires may be initially caused by human activity, but the spread of wildfires is mainly determined by environmental factors, for examples, (1) weather conditions such as temperature, wind direction and intensity, and moisture levels; (2) the amount and types of dry vegetation in a local area, and (3) topographic or local terrian conditions, which affects how much rain an area gets and how fire dynamics will be constrained or faciliated. Thus, to accurately forecast wildfire occurrence has become one of most urgent and taunting environmental challenges in global scale. In this work, we developed a real-time Multimodal Transformer Neural Network Machine Learning model that combines several advanced artificial intelligence techniques and statistical methods to practically forecast the occurrence of wildfire at the precise location in real time, which not only utilizes large scale data information such as hourly weather forecasting data, but also takes into account small scale topographical data such as local terrain condition and local vegetation conditions collecting from Google Earth images to determine the probabilities of wildfire occurrence location at small scale as well as their timing synchronized with weather forecast information. By using the wildfire data in the United States from 1992 to 2015 to train the multimodal transformer neural network, it can predict the probabilities of wildfire occurrence according to the real-time weather forecast and the synchronized Google Earth image data to provide the wildfire occurrence probability in any small location ($100m^{2}$) within 24 hours ahead.
journal: Computer Methods in Applied Mechanics and Engineering
@@ 1 Introduction
Due to climate change, extreme wildfire has become one of the most frequently-happening, widely spreading, and most dangerous hazard on Earth. For example, of nearly \$150 billion loss in direct damages estimated have been caused by California’s 2018 wildfire alone. In addition, wildfire smoke also causes significant environmental pollution. According to an estimate from a report [1], Global forest fires emitted 5 to 8 billion tonnes of $CO_{2}$ each year, and it is more than 10% of the global energy-related $CO_{2}$ emission. In particular, in the recent 2018 California Camp Fire (see Figure 1 (a)), there were 85 confirmed deaths, 17 injuries, and more than $52,000$ people were evacuated. In that fire, more than 18,804 building structures were destroyed, and the burn area exceeds $621~{}km^{2}$. The Camp Fire caused more than \$ 16.65 billion dollars in direct damage alone (2018 USD)). Figure 1 (b) shows the geological distribution of the wildfires occurred in the United States.
@@
Figure 1: (a) 2018 California Camp fire [Photo: U.S. Department of Agriculture/Wikimedia Commons], and (b) Geographical locations of wildfire occurred in the United States
@@
By accurately forecasting the wildfire occurrence location and its timing, we cannot only reduce the indiscriminate power shutdown in the fire season, which causes power outage resulting of closing business and schools, but also allow us take active and specific and direct preventive measures in surgical precision. For instance, if we know the precise fine scale fire occurrence location, we can control the vegetation growth and state in the specific predicted location in a relatively short time with low-cost and high efficiency. Since 2010s, a number of researches have used artificial intelligence-related or machine learning (ML) based methods to predict wildfire occurrences in a relatively large area e.g. [2, 3]. However, most of these research works are still at a scholarly research level but not at the practical fire-hazard prevention stage, such as forecast the precise occurrence location within $1$ mile radial distance. The current ML-based wildfire prediction models used the existing wildfire data and large scale weather data to predict possible fire location. Even though the existing machine learning neural network models may be able to predict human-caused fire occurrence, the predicted location is inside a large area $(>10~{}mile^{2})$. However, the actual and precise wildfire occurrence location in a landscape is strongly influenced by local weather, terrain, and vegetation conditions. Precisely speaking, the precise fire occurrence location depends on the small scale weather, terrain, and vegetation conditions that cannot be obtained from the existing database or weather forecasting, but can only be extrapolated or correlated with the local terrain and vegetation conditions, and some of them can only be represented by the spatial conditional probability distribution. Moreover, in some advanced studies on machine learning-based wildfire occurrence predictions, most researchers adopt the logistic regression statistical analysis (correlation) and artificial neural network (ANN) approach to directly predict fire occurrence by adding vegetation cover[4], slope, elevation, the density of livestock, precipitation conditions, and lightning polarity as features (See: [5, 6, 7]).However, the spatial scale of these features are either inconsistent with the scale of weather forecasts or incompatible with the length scale of the existing wildfire data. Since wildfire occurrence and its severity in a landscape is strongly influenced by both weather and terrain conditions [8] at different locations, and in different spatial and temporal scales, By categorizing the different causes of wildfire, and using and learning from the vast wildfire data, in this work we are developing a multiscale multimodal learning neural network model that can provide the conditional probability of the wildfire for different causes, Unlike the current straightforward and elementary machine learning approach, the proposed Machine Learning Neural Network approach can offer an efficient tool to practically forecast the probability of wildfire occurrence in a small scale spatial location and a precise time frame. It not only takes into account large scale regular weather and historic wildfire data such as hourly weather forecasting data, but also utilizes small scale geographic information such as local terrain conditions and local vegetation conditions extrapolated from Google Earth images to obtain the probabilities of wildfire occurrence with respect to small scale location as well as their timing.
@@ 2 Data Collections and Data Pre-processing
In this work, we only use the wildfire data collected in the United States to demonstrate the feasibility of the developed machine learning neural network. As shown in Figure 2, the wildfire data-set is mainly consisted of two parts. The first part is the information data, and the second part is the imagery data. Within the first part, information data is a combination of three categories or sets of data: wildfire information set, weather information set, and vegetation information set. The first data set, wildfire information, consists of the 24 years, from 1992 to 2015, of geo-referenced wildfire records from the United States Department of Agriculture (USDA) [9]. From the geo-referenced records, it mainly includes the fire location, which is latitude and longitude, the size of the fire, which is treated as quantitative data, and the causes of fire including the date of wildfire was discovered, which are also qualitative data. The qualitative data will be converted into quantitative data by using methods such as One-Hot Encoding [10]. The miscellaneous data such as the fire name, fire ID are not included in this class of data. The second data set, weather information, contains mainly Integrated Surface Data (ISD). It contains the records of the weather conditions before the wildfire occurred. For instance, it may contain the 7 days hourly average temperature before before the discover date from wildfire records. The vegetation information will contain the categories of the location that wildfire was occurred from wildfire records. The second part is the imagery set that we manually obtained by software such as Google Earth Pro[11]. It contains the obtainable images at the specific time when fire occurred within a certain radius from the fire locations as center with a clear view of the surface situations.
@@
Since some data may not be available, when combining two parts, we remove the data with missing values.
@@ 2.2 Imagery Data
The second part of data is the images of the locations, and this part of date contains all the known wildfire location images that were captured by Google Earth Pro[11]. Google Earth Pro provides satellite images for each location, however, these images are not captured continuously. For instance, the time when some images taken in the same location may have several month time difference. Moreover, the obtainable images may not be the exact date when fire occurred. Within the data set, the image adopted will be the first available image before the wildfire discovery date ($disc\_date\_pre$). Moreover, the image quality are also different due to different cameras in different satellites or other causes like sun lights or cloud conditions. These images were taken by the cameras or sensors from multiple satellites such as Landsat, NASA satellite [15], or Sentinel satellites from ESA[16] at different times with different sensor technologies. For example, the most recent United States Geological Survey (USGS) Landsat 9 has ultra blue (coastal aerosol) surface reflectance for Band 1, which allows for coastal water color observation, and Band 2, 3, 4 surface reflectance are blue, green, and red. However, some images were taken by using USGS Landsat 7 that only have blue, green, and red band. Therefore, the quality may varies due to the improvement of satellite and sensor technologies. Furthermore, since the data were taken in a time span of 24 years, most images obtained in 1990s are black-white gray scale images. This can be seen from image number 5536 in Figure 4, and some images may be also in poor quality due to the glare or sunshine, such as the number 11191 image taken from the satellite as shown in Figure 4. But after 2005, with the advance in satellite image technologies, the image quality are much better than those image taken before. Table 1 shows some basic settings when we captured images from Google Earth Pro. Turning setting sun off will result a mask alleviate the effects of sun, which can also help keeping the consistence of the image quality. In Google Earth Engine[11], it automatically masked the clouds if the cloud is in the satellite image, thus, we can obtain ground image without clouds by unchecking the cloud in the settings. The map scale is in default for Google Earth Pro, when we process the images we will crop and re-scale the size of the images. We also removed the road or any other irrelevant marks or pins when capture the images. Thus, as we can see from Figure 4, the images taken from Google Earth Pro is purely terrain without any pins or marks.
@@
To efficiently generate data, we used an automated software to generate the image data. With the help of Pulover’s Marco Creator[17], a software that can automate simple mouse clicks and complex macros with loops and conditions[17], computer is able to automatically captured the images from Google Earth Pro. This software is able to record the action that one makes, then repeats the process until a certain amount images are obtained. In this work, we installed the software and Googel Earth Pro in Dell Precision Tower 5820, and we collected all image data in almost three months. After we gathering the images, we cropped and scaled the images into the size or scale that we desired, which is H: $100\ pixels\times\text{W: }100\ pixels$ within the approximate length of 100 meters by 100 meters where the center is the coordinate of the wildfire occurred.
@@
To obtain 1.88 million images from Google Earth is time consuming even using an automation software. Each image entry may take more than half minutes to finish. not including time delay caused by other reasons such as data requesting and data fetching. Sometimes, some locations do not have satellite images, which will lead the computer pause because of the incorrect date format due to the software default settings.
@@ 2.3 Data Pre-processing
In order to maintain the consistence of image data information and avoid poor contrast, when handling the image data, we transformed and normalized [18] all the images to gray-scaled to keep consistency with the low contrast images, and we normalized the image data before training them by dividing the gray-scaled image data by 255. After the process, we concatenate these two data-sets by the accordance index and generate a new data set including the weather data, vegetation, and processed satellite images from map. Therefore, the ready-to-train data-set has a dimension of 29550 rows and 21 columns plus the image features. Before training, we split the data into three subsets, the first one is temperature and wind, then the second subset includes two more features: humidity and precipitation. In the third subset, we include vegetation category to evaluate the performance of the baseline models with different features. Our current target is to predict the probability of whether a certain nature condition will caused wildfire at certain location, for instance lightning, as a main ignition source of of natural wildfires [19]. Even though some wildfire machine learning models in the literature also claim to be able to predict human-caused fire occurrence, e.g. [20], predicting human caused wildfire is not the objective in this work. Other causes shown in 3. General Wildfire Causes defined in National Wildfire Coordinating Group are also directly or partially directly caused by humanities, for now it is hard to predict them because they are based on human behaviors [21], whose statistical information is not clear. While the cause of power line may not be caused by human activities[22], thus in the data description we can see that the relative frequency of wildfire cause by power line has a extremely small amount comparing to that of other causes, as we can see from Figure 6 (a). For this perspective, as we can see from Figure 10, we can treat this problem as a classification problem. Thus, for predicting the wildfire caused by nature, we can use lightning as label. However, the proportion of the label as lightning has about $4717$ cases among $29550$, as shown in Figure 6 (a), which is only about $15.96\%$ of the entire data. We can also see from Figure 6 (b) that the data points are much more sparse compare to the entire data point in Figure 1 (b), which may lead the data imbalanced [23]. Since about $84\%$ of the binary label is false, and it will have significant effect on the prediction results, which the model will tends to have more false in the results instead of true, which can lead a low true positive rate, because identifying to all false will lead an average $84\%$ accuracy. Thus, we applied a re-sampling method to resolve this issue, which will be discussed in the subsection follows.
@@ 2.3.1 Data Resampling
There are several data re-sampling methods, e.g. undersampling[23], oversampling or SMOTE[24]. In the proposed model, we adopted two method: One is Synthetic Minority, an oversampling technique (SMOTE)[24], that increases minority by analyzing the minority population and generate new samples to increase the amount instead of simply copying the sample to data, which may cause overfit problem (see 5. Synthetic Minority Over-sampling Technique(SMOTE)). Moreover, when including the imagery data, SMOTE will need to generate an amount of images, which may take a long time and computer resources. The other method is undersampling[23] method. For this method, it simply decrease the proportions of the majorities by deleting occurrences of the more frequent class. It can balance uneven dataset by keeping all of the data in the minority class and decreasing the size of the majority class[23]. Nevertheless, it may decrease the train data size. In the present work, we apply this method by randomly choosing a certain amount of data from majorities and combine them with the minority to adjust the proportion. We will put the extra majorities to the test case and increase the size of test case. This will lead a large amount of test cases, however, it actually helps us checking the performance of the model, since most of the data are not being trained, and test data are even large than train data. Therefore, between these two methods, we favor the undersampling approach. Thus, we re-sample the amount of the minority data to a certain level, say, the odds can be $1:1.8$, which is about $35.72\%$ as shown in Figure 5.
Figure 5: Data split detail, Total 29550, natural caused 4717, non-natural caused 8490, train size 10565, test size 2642, total test 18985
Therefore, we randomly choose data from non-natural causes, then combine with $4717$ nature-caused wildfire data to keep the proportion as $35.72\%$. If we set to $1:1$ which is $50\%$ of the proportion will lead a very small train data set with $7547$ data trained and a even larger test set. Thus, we prefer an optimal solution on setting the proportion. Through this method, the test set will be relatively large compare even to the train set, it can help us test the model performance better.
@@
Before training, the data will be first split into two sets: training set and test set. If we adopt the proportion: $80\%$ of training set and $20\%$ test set, it will results a total $10565$ in train set. Based on this setting, the rest of the non-natural causes data will all be in the test set which leads $2642+16343=18985$ test cases. For the purpose of training, small size of training data set may be not good, however a larger test size can provide a more plausible results.

### v2

@@ Abstract
Due to climate change, extreme wildfire has become one of the most dangerous natural hazards to human civilization. Even though, some wildfires may be initially caused by human activity, the spread of wildfires is mainly determined by environmental factors, for example, (1) weather conditions such as temperature, wind direction and intensity, and moisture levels; (2) the amount and types of dry vegetation in a local area, and (3) topographic or local terrain conditions, which affect how much rain an area gets and how fire dynamics will be constrained or facilitated. Thus, to accurately forecast wildfire occurrence has become one of the most urgent and taunting environmental challenges on a global scale. In this work, we developed a real-time Multimodal Transformer Neural Network Machine Learning model that combines several advanced artificial intelligence techniques and statistical methods to practically forecast the occurrence of wildfire at the precise location in real-time, which not only utilizes large-scale data information such as hourly weather forecasting data, but also takes into account small-scale topographical data such as local terrain condition and local vegetation conditions collecting from Google Earth images to determine the probabilities of wildfire occurrence location at small scale as well as their timing synchronized with weather forecast information. By using the wildfire data in the United States from 1992 to 2015 to train the multimodal transformer neural network, it can predict the probabilities of wildfire occurrence according to the real-time weather forecast and the synchronized Google Earth image data to provide the wildfire occurrence probability in any small location ($100m^{2}$) within 24 hours ahead.
@@ 1 Introduction
Due to climate change, extreme wildfire has become one of the most frequently happening, widely spreading, and most dangerous hazards on Earth. For example, the direct property damage from the 2025 wildfire in Los Angeles County is estimated at between \$28 billion and \$53.8 billion, according to a report commissioned by the Southern California Leadership Council. Nearly \$150 billion loss in direct damages has been caused by California’s 2018 wildfire alone. In addition, wildfire smoke also causes significant environmental pollution. According to an estimate from a report [1], Global forest fires emitted 5 to 8 billion tonnes of $CO_{2}$ each year, and it is more than 10% of the global energy-related $CO_{2}$ emission. In particular, the 2025 wildfire in Los Angeles County have killed at least 29 people, and forced more than 200,000 people to evacuate. While in the 2018 California Camp Fire (see Figure 1 (a)), there were 85 confirmed deaths, 17 injuries, and more than $52,000$ people were evacuated. In that fire, more than 18,804 building structures were destroyed, and the burn area exceeded $621~{}km^{2}$. The Camp Fire caused more than 16.65 billion dollars in direct damage alone (2018 USD)). Figure 1 (b) shows the geological distribution of the wildfire that occurred in the United States.
@@
Figure 1: (a) 2018 California Camp Fire [Photo: U.S. Department of Agriculture/Wikimedia Commons], and (b) Geographical locations of wildfires occurred in the United States
@@
By accurately forecasting the wildfire occurrence location and its timing, we cannot only reduce the indiscriminate power shutdown in the fire season, which causes power outages resulting of closing businesses and schools, but also allow us to take active specific and direct preventive measures in surgical precision. For instance, if we know the precise fine-scale fire occurrence location, we can control the vegetation growth and state in the specific predicted location in a relatively short time with low cost and high efficiency. Since the 2010s, several studies have used artificial intelligence-related or machine learning (ML) based methods to predict wildfire occurrences in a relatively large area e.g. [2, 3]. However, most of these research works are still at a scholarly research level but not at the practical fire-hazard prevention stage, such as forecast the precise occurrence location within $1$ mile radial distance. The current ML-based wildfire prediction models used the existing wildfire data and large-scale weather data to predict possible fire locations. Even though the existing machine-learning neural network models may be able to predict some fire occurrences, the predicted location is inside a large area $(>10~{}mile^{2})$. However, the actual and precise wildfire occurrence location in a landscape is strongly influenced by local weather, terrain, and vegetation conditions. Precisely speaking, the precise fire occurrence location depends on the small-scale weather, terrain, and vegetation conditions that cannot be obtained from the existing database or weather forecasting, but can only be extrapolated or correlated with the local terrain and vegetation conditions, and some of them can only be represented by the spatial conditional probability distribution. Moreover, in some advanced studies on machine learning-based wildfire occurrence predictions, most researchers adopt the logistic regression statistical analysis (correlation) and artificial neural network (ANN) approach to directly predict fire occurrence by adding vegetation cover[4], slope, elevation, the density of livestock, precipitation conditions, and lightning polarity as features (See: [5, 6, 7]).However, the spatial scale of these features is either inconsistent with the scale of weather forecasts or incompatible with the length scale of the existing wildfire data. Since wildfire occurrence and its severity in a landscape is strongly influenced by both weather and terrain conditions [8] at different locations, and in different spatial and temporal scales, By categorizing the different causes of wildfire, and using and learning from the vast wildfire data, in this work, we are developing a multiscale multimodal learning neural network model that can provide the conditional probability of the wildfire for different causes, Unlike the current straightforward and elementary machine learning approach, the proposed Machine Learning Neural Network approach can offer an efficient tool to practically forecast the probability of wildfire occurrence in a small-scale spatial location and a precise time frame. It not only takes into account large-scale regular weather and historic wildfire data such as hourly weather forecasting data, but also utilizes small-scale geographic information such as local terrain conditions and local vegetation conditions extrapolated from Google Earth images to obtain the probabilities of wildfire occurrence with respect to small-scale location as well as their timing.
@@ 2 Data Collections and Data Pre-processing
In this work, the wildfire data collected are only within the United States to demonstrate the feasibility of the developed machine learning neural network. As shown in Figure 2, the wildfire data is mainly consists of two parts. The first part is the information data, and the second part is the imagery data. Within the first part, information data is a combination of three categories or sets of data: wildfire information set, weather information set, and vegetation information set. The first data set is wildfire information, which consists of geo-referenced wildfire records from the United States Department of Agriculture (USDA) [9] from 1992 to 2015 of 24 years, The geo-referenced records mainly include the fire location, which is latitude and longitude, the size of the fire, which is treated as quantitative data, and the causes of fire including the date of the wildfire, which is also qualitative data. In this study, the qualitative data will be converted into quantitative data by using methods such as One-Hot Encoding [10]. The miscellaneous data such as the fire name, and the fire ID are not included in this class of data. The second data set, weather information, contains mainly Integrated Surface Data (ISD). It contains the records of the weather conditions before the wildfire occurred. For instance, it may contain the 7 days hourly average temperature before the discovery date in the wildfire records. The vegetation information will contain the categories of the location that wildfire was occurred in the wildfire records. The second part is the imagery set that is manually obtained by software such as Google Earth Pro[11]. It contains the obtainable images at the specific time when the fire occurred within a certain radius from the fire locations as the center with a clear view of the surface situations.
@@
Since some data may not be available when combining two parts, we remove the data with missing values.
@@ 2.2 Imagery Data
The second part of the data are the images of the locations, and this part of the data contains all the known wildfire location images that was captured by Google Earth Pro[11]. Google Earth Pro provides satellite images for each location, however, these images are not captured continuously. For instance, the time when some images are taken in the same location may have a several-month time difference. Moreover, the obtainable images may not be the exact date when a fire occurred. Within the data set, the image adopted will be the first available image before the wildfire discovery date ($disc\_date\_pre$). Moreover, the image qualities in different pictures are also different due to the different cameras on different satellites used taking those pictures, or due to other causes such as different sunlight or cloud conditions. These images were taken by the cameras or sensors from multiple satellites such as Landsat, NASA satellite [15], or Sentinel satellites from ESA[16] at different times with different sensor technologies. For example, the most recent United States Geological Survey (USGS) Landsat 9 has ultra blue (coastal aerosol) surface reflectance for Band 1, which allows for coastal watercolor observation, and Band 2, 3, and 4 surface reflectance are blue, green, and red. However, some images were taken by using USGS Landsat 7 that only have blue, green, and red bands. Therefore, the quality may vary due to the improvement of satellite and sensor technologies. Furthermore, since the data were taken in 24 years, most images obtained in the 1990s are black-white grayscale images. This can be seen from image number 5536 in Figure 4, and some images may be also in poor quality due to the glare or sunshine, such as the number 11191 image taken from the satellite as shown in Figure 4. But after 2005, with the advance in satellite image technologies, the image quality is much better than those images taken before. Table 1 shows some basic settings when the images are captured from Google Earth Pro. Turning the setting sun off will result in a mask alleviating the effects of the sun, which can also help keep the consistency of the image quality. In Google Earth Engine[11], it automatically masks the clouds if the cloud is in the satellite image, thus, one can obtain ground images without clouds by unchecking the cloud in the settings. The map scale is in default for Google Earth Pro, when the images are being processed and it will crop and re-scale the size of the images. The road or any other irrelevant marks or pins are also removed when capturing the images. Thus, as shown in Figure 4, the images taken from Google Earth Pro are purely terrain-type without any pins or marks.
@@
To efficiently generate data, automated software is used to generate the image data. With the help of Pulover’s Marco Creator[17], a software that can automate simple mouse clicks and complex macros with loops and conditions[17] is used, which can automatically capture the images from Google Earth Pro. This software can record the action that one makes, then repeat the process until a certain amount of images are obtained. In this work, we installed the software and Google Earth Pro in a Dell Precision Tower 5820 desktop, and we collected all image data in almost three months. After we gathered the images, we cropped and scaled the images into the size or scale that desired, which is H: $100\ pixels\times\text{W: }100\ pixels$ within the approximate length of 100 meters by 100 meters where the center is the coordinate of the wildfire occurred.
@@
To obtain 1.88 million images from Google Earth is time-consuming even using automation software. Each image entry may take more than half a minute to finish, not including time delay caused by other reasons such as data requesting and data fetching. Sometimes, some locations do not have satellite images, which will lead the computer to pause because of the incorrect date format due to the software’s default settings.
@@ 2.3 Data Pre-processing
To maintain the consistency of image data information and avoid poor contrast, when handling the image data, the data were transformed and normalized [18] all the images to gray-scaled to keep consistency with the low contrast images, and it will be normalized the image data before training them by dividing the gray-scaled image data by 255. After the process, these two data sets are concatenated by the accordance index and generate a new data set including the weather data, vegetation, and processed satellite images from the map. Therefore, the ready-to-train data set has a dimension of 29550 rows and 21 columns plus the image features. Before training, the data were split into three subsets, the first one is temperature and wind, then the second subset includes two more features: humidity and precipitation. In the third subset, vegetation categories are included to evaluate the performance of the baseline models with different features. Our current target is to predict the probability of whether a certain natural condition will cause wildfire at certain locations, for instance lightning, as a main ignition source of natural wildfire [19]. Even though some wildfire machine learning models in the literature also claim to be able to predict human-caused fire occurrence, e.g. [20], predicting human-caused wildfire is not the objective of the present work. Other causes shown in 3. General Wildfire Causes defined in National Wildfire Coordinating Group are also directly or partially directly caused by humanities, for now, it is hard to predict them because they are based on human behaviors [21], whose statistical information is not clear. While the cause of power line may not be caused by human activities[22], thus in the data description one can see that the relative frequency of wildfire caused by power lines has an extremely small amount compared to that of other causes, as shown from Figure 6 (a). For this perspective, as shown in Figure 10, one can treat this problem as a classification problem. Thus, for predicting wildfires caused by nature, lightning can be used as a label. However, the proportion of the label as lightning has about $4717$ cases among $29550$, as shown in Figure 6 (a), which is only about $15.96\%$ of the entire data. From Figure 6 (b) that the data points are much more sparse compared to the entire data point in Figure 1 (b), which may lead the data imbalanced [23]. Since about $84\%$ of the binary label is false, and it will have a significant effect on the prediction results, which the model will tend to have more false in results instead of true, which can lead to a low true positive rate, because identifying all false will lead to an average $84\%$ accuracy. Thus, a re-sampling method could resolve this issue, which will be discussed in the subsection follows.
@@ 2.3.1 Data Resampling
There are several data re-sampling methods, e.g. undersampling[23], oversampling or SMOTE[24]. In the proposed model, two methods are applied: One is Synthetic Minority, an oversampling technique (SMOTE)[24], that increases minority by analyzing the minority population and generate new samples to increase the amount instead of simply copying the sample to data, which may cause an overfitting problem (see 5. Synthetic Minority Over-sampling Technique(SMOTE)). The other method used is the undersampling[23] method. This method simply decreases the proportions of the majorities by deleting occurrences of the more frequent class. It can balance uneven datasets by keeping all of the data in the minority class and decreasing the size of the majority class[23]. Nevertheless, it may decrease the train data size. In the present work, this method is used by randomly choosing a certain amount of data from majorities and combining them with the minority to adjust the proportion. By putting the extra majorities to the test case, these methods increase the size of the test cases. However, it helps us check the performance of the model, since most of the data are not being trained, and test data are even larger than train data. Therefore, between these two methods, we favor the undersampling approach. Thus, re-sampling the amount of the minority data to a certain level, say, the odds can be $1:1.8$, which is about $35.72\%$ as shown in Figure 5 is an optimal choice. Therefore, we randomly choose data from non-natural causes, then combine it with $4717$ nature-caused wildfire data to keep the proportion as $35.72\%$. If we set the proportion to $1:1$, which is $50\%$ for the training and the test data, it will lead to a very small training data set with $7547$ data trained and a larger test set. Thus, we prefer an optimal solution of setting the proportion. Through this method, the test set will be relatively large compared even to the train set, it can help us test the model performance better.
Figure 5: Data split detail, Total 29550, naturally caused 4717, non-natural caused 8490, train size 10565, test size 2642, total test 18985
Nevertheless, this may decrease the training data size. In the present work, we apply the method by randomly choosing a certain amount of data from majorities and combine them with the minority to adjust the proportion. We will put the extra majorities to the test case and increase the size of test case. This will lead a large amount of test cases, however, it actually helps us checking the performance of the model, since most of the data are not being trained, and test data are even large than train data. Therefore, between these two methods, we favor the undersampling approach. Thus, we re-sample the amount of the minority data to a certain level, say, the odds can be $1:1.8$, which is about $35.72\%$ as shown in Figure 5.
@@
Before training, the data will be first split into two sets: the training set and the test set. If the proportion of the total data is about $80\%$ of the training set and $20\%$ test set, it will have a total of $10565$ in the training set. Based on this setting, the rest of the non-natural causes data will all be in the test set which leads to $2642+16343=18985$ test cases. For the training purpose alone, a small size training data set may not be good, however, a larger test size can provide more plausible results. This method also helps to check whether the model is overfitted during training. Most of the non-natural causes are not included in the model during training using undersampling. When testing the model, these cases are the model never seen during training. Thus, if the model still performs well during testing, the model is not affected by the proportion, and it will not be overfitted.